# Professor Leonardo P. Chamorro Professor Cristian Ricardo Constante Amores Professor Jie Feng

### v1

@@ 2.1 Introduction
The motion of particles suspended in viscous flows near boundaries is a fundamental problem in low Reynolds number hydrodynamics with applications in microfluidics, colloidal suspensions, and biomedical systems. In most microfluidic setups, at least a subset of particles placed into the flow will be located or transported near boundaries, which could be solid (no-slip) or fluid/fluid interfaces (e.g. immiscible liquids, droplets, bubbles). Generally, the proximity of a boundary leads to specific displacements, depending on the Reynolds number, density contrast, and other parameters. In steady inertial microfluidics the presence of boundaries and the resulting flow gradients lead to slow inertial migration even at considerable distances to the boundaries (segre1962behaviour, ; di2009inertial, ), while in oscillatory inertial microfluidics (such as setups using acoustically driven microbubbles), boundary effects become important in very close proximity and can be approximately treated by lubrication theory (thameem2017fast, ; agarwal2018inertial, ). In a flow with negligible inertia, one would expect boundary effects to be longer-range and potentially more prominent. Still, the question of whether a practically usable net displacement after an encounter of a particle with a wall or an obstacle is feasible has not been fundamentally answered. Although the motion of a force-free particle in an unbounded fluid is well understood, the effects of proximity to rigid walls are still lacking. The ability to quantify these wall-induced effects is critical for predicting transport, mixing, and sorting behaviors in confined systems.
Different quantities can be targeted in the modeling of particles in Stokes flow, particularly (i) the forces on a particle moving at a given speed, (ii) the forces on a particle held fixed in a certain location, or (iii) the motion of a force-free particle. The latter is our focus here, as it describes the trajectory of a density-matched particle that is not subject to external forces. Progress in describing all three cases has been built on a body of literature based on early pioneering work (brenner1961slow, ; goldman1967slow, ; goldman1967slow2, ).
Generally, studies of force-free particles result in predictions for the deviation of the particle velocity from the background fluid velocity in which it is embedded, i.e., the non-passive part of the particle motion. This velocity correction can be decomposed into effects in the direction parallel to the boundary and in the perpendicular direction. Previous analysis of the equation of motion of a force-free particle with wall-normal velocity corrections was given by (rallabandi2017hydrodynamic, ). Other work has described wall-parallel velocity corrections far from and near the wall (ekiel2006accuracy, ; pasol2011motion, ).
@@ 2.2 Moffatt Eddies
The ideal test case to quantify boundary effects in a Stokes flow is a flow that is (i) analytically known and for which (ii) the walls are isolated and flat. We take inspiration from the classic work of (moffatt1964viscous, ): when two rigid flat boundaries form a wedge, a distant stirring of the fluid will induce a flow consisting of a sequence of vortices shown in Fig. 2.2 (a). (moffatt1964viscous, ) also describes the special case of zero wedge angle, i.e., a Stokes flow between two parallel plates as sketched in Fig. 2.2 (b). We set these two plates at $y=\pm 1$.
Figure 2.2: (a) Schematic of Stokes flow in a wedge between rigid boundaries. (b) Schematic of Stokes flow between parallel plates; the source of the fluid motion can be visualized as a rotating cylinder at a large negative $x$ coordinate between the plates. Modified from the original sketches in (moffatt1964viscous, ).
@@
The constant $A$ is an overall scale that can be set to 1. In order to fulfill the no-slip boundary conditions at $y=\pm 1$, $B=-\cot(k_{S})$ follows. Furthermore, the complex parameter $k_{S}=p_{S}+iq_{S}$ must satisfy the transcendental equation $2k_{S}+\sin 2k_{S}=0$ (moffatt1964viscous, ). The solution with the smallest positive real part is $p_{S}\approx 2.106,q_{S}\approx 1.125$. Its symmetry with respect to the center plane of the channel implies (together with the time-reversal symmetry of Stokes flow) that a particle released anywhere and deflected by interaction with one of the walls will experience the opposite deflection when encountering the other wall so that all trajectories must close and there is no net displacement of particles.
@@ 2.3 Particle velocity in the presence of a wall
In any ambient Stokes flow, the velocity of a spherical particle of radius $a_{p}$ differs from the background flow velocity $\bf{u}$ (without walls) by the Faxen correction (faxen1922widerstand, ) evaluated at the position of the particle $\bf{x}_{p}$, resulting in
@@
We will show that in many situations for small $a_{p}$ the wall interaction $\bm{W}$ is perturbative, i.e., of a higher order than $a_{p}^{2}$. The principle of (2.5) has been acknowledged in the literature (brenner1961slow, ; goldman1967slow, ; goldman1967slow2, ; o1964slow, ; o1967slow, ; o1967slow2, ; perkins1992hydrodynamic, ), but has not been systematically applied for arbitrary $h$ to determine particle trajectories meant for net displacement. The following subsections quantify the particle velocity corrections $\bm{W}$ parallel to and normal to the walls.
@@ 2.4 Wall-parallel corrections to the particle velocity
representing the surface-to-surface distance relative to the radius of the particle, cf. (rallabandi2017hydrodynamic, ; thameem2017fast, ; agarwal2018inertial, ).
The wall-parallel velocity correction coefficient $f(\Delta)$ has been worked out in detail for specific cases such as linear shear flow (o1968sphere, ; jeffrey1984calculation, ; stephen1992characterization, ; williams1994particle, ; chaoui2003creeping, ), quadratic flow (goren1971hydrodynamic, ; ekiel2006accuracy, ; pasol2006sphere, ), or modulated shear flow (pasol2006sphere, ), with asymptotic expressions available for $\Delta\to 0$ and $\Delta\to\infty$. We note that (i) the wall interactions are most prominent for small $\Delta$ and (ii) the linear shear part of any flow dominates as $\Delta\to 0$. In particular, when $\Delta=0$ (particle touching the wall), the sphere has to come to rest.
In order to obtain a uniformly valid expression for $f(\Delta)$, we follow the expansion approach of (pasol2011motion, ) but modify it to enforce exact matching with known asymptotic results. For $\Delta\ll 1$, linear shear flow is dominant, and Williams’s near-wall expression (williams1994particle, ) must be recovered. Far from the wall, $f(\Delta)\to c\Delta^{-3}$, where the positive constant $c={\cal O}(1)$ depends on the type of flow (goldman1967slow2, ; ghalia2016sphere, ). The exact value of $c$ makes no qualitative difference to the effects explored here, and we enforce $c=5/16$ to agree with the far-field asymptote provided by (goldman1967slow2, ) for linear shear flow.
@@
employed for all $\Delta$. Fig. 2.4(b,c) illustrate the agreement with the asymptote at $\Delta\gg 1$ (goldman1967slow2, ) as well as the logarithmic lubrication theory approach to $f=1$ at $\Delta\to 0$ (stephen1992characterization, ; williams1994particle, ).
@@ 2.5 Wall-normal corrections to the particle velocity
A general expression for the wall-normal component of the hydrodynamic force on spherical particles in arbitrary Stokes flows was obtained by (rallabandi2017hydrodynamic, ), employing a quadratic expansion of the background flow around the center of the particle:
@@
where $\bm{e_{\perp}}$ is the unit normal to the wall pointing toward the particle center, and $\mu$ is the viscosity of the fluid. $\bm{u}_{p}=(u_{p},v_{p})$ denotes the velocity of the particle when employing an expansion around the particle position. The scalar quantities $\mathcal{A}$, $\mathcal{B}$, $\mathcal{C}$ and $\mathcal{D}$ are analytically known dimensionless hydrodynamic resistances depending on $\Delta$. The first correction term, proportional to $\mathcal{A}$, is due to the translation of the particle relative to the mean surrounding background flow and is identical to the general expression from (brenner1961slow, ). The term proportional to $\mathcal{B}$ is due to the extensional gradients of the flow field. This contribution is zero for a sphere in an infinite flow field (happel1965low, ; batchelor1970stress, ) but is generally nonzero for a finite distance to the wall. The second moments of the background flow result in two separate contributions to the force: one (proportional to $\mathcal{C}$) dependent on the curvature of the background flow velocity normal to the wall ($\bm{e_{\perp}}\bm{e_{\perp}}:\nabla\nabla\bm{u}$); another (proportional to $\mathcal{D}$) proportional to $\nabla^{2}\bm{u}$. This latter term asymptotes to the Faxen correction as $\Delta\to\infty$.
The full analytical expressions for $\mathcal{A}$, $\mathcal{B}$, $\mathcal{C}$ and $\mathcal{D}$ are given in (rallabandi2017hydrodynamic, ). The asymptotic behaviors of $\mathcal{A},\mathcal{B},\mathcal{C},\mathcal{D}$ for large separations $\Delta\gg 1$ ($\mathcal{A}_{large}$ etc.) and for small separations $\Delta\ll 1$ ($\mathcal{A}_{small}$ etc.) are given in Appendix B.
@@
and it is understood that $\mathcal{A},\mathcal{B},\mathcal{C},\mathcal{D}$ are evaluated at arguments $\Delta$ from (2.7) defined by $h=1\pm y_{p}$ for the walls at $y=\mp 1$, respectively. Note that the last term of (2.11) explicitly subtracts the Faxen correction so that each $W_{y}$ term vanishes as $\Delta\to\infty$. When the particle approaches a wall closely, we enter the regime of $\Delta\ll 1$. In the wall-parallel direction, we still use the expression for $W_{x}$ from equation (2.6), which includes the lubrication limit at very small $\Delta$. However, in the wall-normal direction, the formalism relying on Taylor expansion of the flow around the particle center $v_{p}^{PE}$ is not accurate enough to describe the particle motion – this is easily seen because the predicted particle normal velocity from (2.10) does not vanish when the particle touches the wall. In (rallabandi2017hydrodynamic, ), it was shown that for $\Delta\ll 1$, the no-penetration boundary condition can be enforced by replacing the background flow field in (2.11) by its quadratic expansion around the point on the wall closest to the particle. However, we find that this wall-expansion formalism will not smoothly transition to the expression (2.11) when $\Delta\sim 1$, because the exponential dependence of the flow field (2.2) with substantial $|k_{A}|$ compromises the accuracy of the quadratic expansion even at relatively short distances from the wall.
@@ 3.1 Introduction
In the context of Stokes flows, Moffatt eddies offer a paradigmatic example that exhibits nested vortical structures near sharp corners or within confined geometries. In their symmetric form, these flows produce nested, self-similar vortices, supporting closed streamlines and conservative particle motion. In the following, we demonstrate that a neutrally buoyant spherical particle in a symmetry-broken Moffatt eddy flow can exhibit systematic (net) displacement across streamlines due to wall interactions, a phenomenon that has received little attention in the prior literature. These behaviors, while reminiscent of simplified steady-streaming effects (Fig. 1.3), are richer and more complex. While particles can never cross streamlines in unidirectional Stokes flow (bretherton1962motion, ), streamline crossing should generally be expected in the presence of wall-normal flow components. A simple example is a particle very close to a wall transported in a channel flow undergoing contraction; the particle cannot stay on its initial streamline without penetrating the wall.
@@ 3.3.1 Linear stability analysis
where the constant $c$ can vary depending on the specific flow but is ${\cal O}(1)$ (goldman1967slow2, ; ghalia2016sphere, ), cf. Sec. 2.4. Near the fixed point of the Faxen field, the $\frac{\mathcal{B}_{large}}{\mathcal{A}_{large}}$ term in $W_{y}$ dominates the others ($\frac{\mathcal{B}_{large}}{\mathcal{A}_{large}}=\mathcal{O}(a_{p}^{3}),\frac{\mathcal{C}_{large}}{\mathcal{A}_{large}}=\mathcal{O}(a_{p}^{5}),\frac{\mathcal{D}_{large}}{\mathcal{A}_{large}}=\mathcal{O}(a_{p}^{4})$) and is the only one contributing to $\mathcal{O}(a_{p}^{3})$ in (3.2). This term is still of higher $a_{p}$ order than the Faxen correction, so that, in the limits of large $\Delta$ and small $a_{p}$, the wall correction is perturbatively small.
@@ 3.5 Particle motion and manipulation in a counterclockwise vortex
Any vortex adjacent to a clockwise vortex, like the one discussed in Sec. 3.3, is counterclockwise and congruent in geometry. For example, the vortex indicated by the green frame in Fig. 2.4(a) has flow exactly reversed from the yellow-framed vortex (and a factor $\zeta$ slower). Because of the time reversibility of Stokes flow and the fact that all wall interactions result from the background Stokes flow and its derivatives, the behavior of particles on trajectories is also time-reversed. Thus, the fixed point in a counterclockwise vortex is stable, $\frac{1}{\tau}$ changes sign, and $\beta$ stays the same by definition. The limit cycle for a given $a_{p}$ is unstable but is congruent in shape with the stable cycle discussed earlier. Particles spiral inward from the unstable limit cycle towards the fixed points, but spiral outward when placed outside the limit cycle (Fig. 3.3(a)). The latter case is of prime interest because it allows particles to approach the wall more and more closely. As the gap between particle and wall diminishes, any short-ranged intermolecular attractive force (e.g. Van der Waals force) whose reach is often in the nm range (hirschfelder1954molecular, ; batsanov2001van, ) can take over and lead to attachment (sticking) of the particle to the wall. This general mechanism (hydrodynamics allows a particle to get close enough to a wall to stick by short-range attraction) has been acknowledged before (friedlander2000smoke, ; humphries2009filter, ), but the case discussed here allows for quantitative predictions.
@@
where $\kappa(x)=\frac{\partial^{2}v}{\partial y^{2}}(x,-1)$ is the background flow curvature at the wall, and the prefactor was first derived in (goren1971hydrodynamic, ) and confirmed in (rallabandi2017hydrodynamic, ). The part of the trajectory showing the exponential approach (the solid portion in Fig. 3.3(b,c)) indicates slow motion in the $x$-direction, so that $\kappa(x)$ is nearly constant. We fix the $x$-value as that of the turning point (point of maximum curvature) of the trajectory, here $x\approx 0.991$. Using this value in (3.11) gives an exponential behavior (dashed line in Fig. 3.3(c)) in good agreement with the approach of the trajectory to the wall. This behavior remains robust to changes in the exact modeling of the expansion point locations $y_{E}$ in (2.13).
@@
Our spherical-particle formalism implicitly assumes the limit of dilute particle concentration in a microfluidics application. Including the effects of particle-particle interaction is a desirable future extension of this approach, allowing for assessment of non-inertial effects in particle-laden flows, which are important for practical applications (guha2008transport, ).

### v2

@@ 2.1 Introduction
The motion of particles suspended in viscous flows near boundaries is a fundamental problem in low Reynolds number hydrodynamics with applications in microfluidics, colloidal suspensions, and biomedical systems. In most microfluidic setups, at least a subset of particles placed into the flow will be located or transported near boundaries, which could be solid (no-slip) or fluid/fluid interfaces (e.g. immiscible liquids, droplets, bubbles). Generally, the proximity of a boundary leads to specific displacements, depending on the Reynolds number, density contrast, and other parameters. In steady inertial microfluidics the presence of boundaries and the resulting flow gradients lead to slow inertial migration even at considerable distances to the boundaries (segre1962behaviour; di2009inertial), while in oscillatory inertial microfluidics (such as setups using acoustically driven microbubbles), boundary effects become important in very close proximity and can be approximately treated by lubrication theory (thameem2017fast; agarwal2018inertial). In a flow with negligible inertia, one would expect boundary effects to be longer-range and potentially more prominent. Still, the question of whether a practically usable net displacement after an encounter of a particle with a wall or an obstacle is feasible has not been fundamentally answered. Although the motion of a force-free particle in an unbounded fluid is well understood, the effects of proximity to rigid walls are still lacking. The ability to quantify these wall-induced effects is critical for predicting transport, mixing, and sorting behaviors in confined systems.
Different quantities can be targeted in the modeling of particles in Stokes flow, particularly (i) the forces on a particle moving at a given speed, (ii) the forces on a particle held fixed in a certain location, or (iii) the motion of a force-free particle. The latter is our focus here, as it describes the trajectory of a density-matched particle that is not subject to external forces. Progress in describing all three cases has been built on a body of literature based on early pioneering work (brenner1961slow; goldman1967slow; goldman1967slow2).
Generally, studies of force-free particles result in predictions for the deviation of the particle velocity from the background fluid velocity in which it is embedded, i.e., the non-passive part of the particle motion. This velocity correction can be decomposed into effects in the direction parallel to the boundary and in the perpendicular direction. Previous analysis of the equation of motion of a force-free particle with wall-normal velocity corrections was given by (rallabandi2017hydrodynamic). Other work has described wall-parallel velocity corrections far from and near the wall (ekiel2006accuracy; pasol2011motion).
@@ 2.2 Moffatt Eddies
The ideal test case to quantify boundary effects in a Stokes flow is a flow that is (i) analytically known and for which (ii) the walls are isolated and flat. We take inspiration from the classic work of (moffatt1964viscous): when two rigid flat boundaries form a wedge, a distant stirring of the fluid will induce a flow consisting of a sequence of vortices shown in Fig. 2.2 (a). (moffatt1964viscous) also describes the special case of zero wedge angle, i.e., a Stokes flow between two parallel plates as sketched in Fig. 2.2 (b). We set these two plates at $y=\pm 1$.
Figure 2.2: (a) Schematic of Stokes flow in a wedge between rigid boundaries. (b) Schematic of Stokes flow between parallel plates; the source of the fluid motion can be visualized as a rotating cylinder at a large negative $x$ coordinate between the plates. Modified from the original sketches in (moffatt1964viscous).
@@
The constant $A$ is an overall scale that can be set to 1. In order to fulfill the no-slip boundary conditions at $y=\pm 1$, $B=-\cot(k_{S})$ follows. Furthermore, the complex parameter $k_{S}=p_{S}+iq_{S}$ must satisfy the transcendental equation $2k_{S}+\sin 2k_{S}=0$ (moffatt1964viscous). The solution with the smallest positive real part is $p_{S}\approx 2.106,q_{S}\approx 1.125$. Its symmetry with respect to the center plane of the channel implies (together with the time-reversal symmetry of Stokes flow) that a particle released anywhere and deflected by interaction with one of the walls will experience the opposite deflection when encountering the other wall so that all trajectories must close and there is no net displacement of particles.
@@ 2.3 Particle velocity in the presence of a wall
In any ambient Stokes flow, the velocity of a spherical particle of radius $a_{p}$ differs from the background flow velocity $\bf{u}$ (without walls) by the Faxen correction (faxen1922widerstand) evaluated at the position of the particle $\bf{x}_{p}$, resulting in
@@
We will show that in many situations for small $a_{p}$ the wall interaction $\bm{W}$ is perturbative, i.e., of a higher order than $a_{p}^{2}$. The principle of (2.5) has been acknowledged in the literature (brenner1961slow; goldman1967slow; goldman1967slow2; o1964slow; o1967slow; o1967slow2; perkins1992hydrodynamic), but has not been systematically applied for arbitrary $h$ to determine particle trajectories meant for net displacement. The following subsections quantify the particle velocity corrections $\bm{W}$ parallel to and normal to the walls.
@@ 2.4 Wall-parallel corrections to the particle velocity
representing the surface-to-surface distance relative to the radius of the particle, cf. (rallabandi2017hydrodynamic; thameem2017fast; agarwal2018inertial).
The wall-parallel velocity correction coefficient $f(\Delta)$ has been worked out in detail for specific cases such as linear shear flow (o1968sphere; jeffrey1984calculation; stephen1992characterization; williams1994particle; chaoui2003creeping), quadratic flow (goren1971hydrodynamic; ekiel2006accuracy; pasol2006sphere), or modulated shear flow (pasol2006sphere), with asymptotic expressions available for $\Delta\to 0$ and $\Delta\to\infty$. We note that (i) the wall interactions are most prominent for small $\Delta$ and (ii) the linear shear part of any flow dominates as $\Delta\to 0$. In particular, when $\Delta=0$ (particle touching the wall), the sphere has to come to rest.
In order to obtain a uniformly valid expression for $f(\Delta)$, we follow the expansion approach of (pasol2011motion) but modify it to enforce exact matching with known asymptotic results. For $\Delta\ll 1$, linear shear flow is dominant, and Williams’s near-wall expression (williams1994particle) must be recovered. Far from the wall, $f(\Delta)\to c\Delta^{-3}$, where the positive constant $c={\cal O}(1)$ depends on the type of flow (goldman1967slow2; ghalia2016sphere). The exact value of $c$ makes no qualitative difference to the effects explored here, and we enforce $c=5/16$ to agree with the far-field asymptote provided by (goldman1967slow2) for linear shear flow.
@@
employed for all $\Delta$. Fig. 2.4(b,c) illustrate the agreement with the asymptote at $\Delta\gg 1$ (goldman1967slow2) as well as the logarithmic lubrication theory approach to $f=1$ at $\Delta\to 0$ (stephen1992characterization; williams1994particle).
@@ 2.5 Wall-normal corrections to the particle velocity
A general expression for the wall-normal component of the hydrodynamic force on spherical particles in arbitrary Stokes flows was obtained by (rallabandi2017hydrodynamic), employing a quadratic expansion of the background flow around the center of the particle:
@@
where $\bm{e_{\perp}}$ is the unit normal to the wall pointing toward the particle center, and $\mu$ is the viscosity of the fluid. $\bm{u}_{p}=(u_{p},v_{p})$ denotes the velocity of the particle when employing an expansion around the particle position. The scalar quantities $\mathcal{A}$, $\mathcal{B}$, $\mathcal{C}$ and $\mathcal{D}$ are analytically known dimensionless hydrodynamic resistances depending on $\Delta$. The first correction term, proportional to $\mathcal{A}$, is due to the translation of the particle relative to the mean surrounding background flow and is identical to the general expression from (brenner1961slow). The term proportional to $\mathcal{B}$ is due to the extensional gradients of the flow field. This contribution is zero for a sphere in an infinite flow field (happel1965low; batchelor1970stress) but is generally nonzero for a finite distance to the wall. The second moments of the background flow result in two separate contributions to the force: one (proportional to $\mathcal{C}$) dependent on the curvature of the background flow velocity normal to the wall ($\bm{e_{\perp}}\bm{e_{\perp}}:\nabla\nabla\bm{u}$); another (proportional to $\mathcal{D}$) proportional to $\nabla^{2}\bm{u}$. This latter term asymptotes to the Faxen correction as $\Delta\to\infty$.
The full analytical expressions for $\mathcal{A}$, $\mathcal{B}$, $\mathcal{C}$ and $\mathcal{D}$ are given in (rallabandi2017hydrodynamic). The asymptotic behaviors of $\mathcal{A},\mathcal{B},\mathcal{C},\mathcal{D}$ for large separations $\Delta\gg 1$ ($\mathcal{A}_{large}$ etc.) and for small separations $\Delta\ll 1$ ($\mathcal{A}_{small}$ etc.) are given in Appendix B.
@@
and it is understood that $\mathcal{A},\mathcal{B},\mathcal{C},\mathcal{D}$ are evaluated at arguments $\Delta$ from (2.7) defined by $h=1\pm y_{p}$ for the walls at $y=\mp 1$, respectively. Note that the last term of (2.11) explicitly subtracts the Faxen correction so that each $W_{y}$ term vanishes as $\Delta\to\infty$. When the particle approaches a wall closely, we enter the regime of $\Delta\ll 1$. In the wall-parallel direction, we still use the expression for $W_{x}$ from equation (2.6), which includes the lubrication limit at very small $\Delta$. However, in the wall-normal direction, the formalism relying on Taylor expansion of the flow around the particle center $v_{p}^{PE}$ is not accurate enough to describe the particle motion – this is easily seen because the predicted particle normal velocity from (2.10) does not vanish when the particle touches the wall. In (rallabandi2017hydrodynamic), it was shown that for $\Delta\ll 1$, the no-penetration boundary condition can be enforced by replacing the background flow field in (2.11) by its quadratic expansion around the point on the wall closest to the particle. However, we find that this wall-expansion formalism will not smoothly transition to the expression (2.11) when $\Delta\sim 1$, because the exponential dependence of the flow field (2.2) with substantial $|k_{A}|$ compromises the accuracy of the quadratic expansion even at relatively short distances from the wall.
@@ 3.1 Introduction
In the context of Stokes flows, Moffatt eddies offer a paradigmatic example that exhibits nested vortical structures near sharp corners or within confined geometries. In their symmetric form, these flows produce nested, self-similar vortices, supporting closed streamlines and conservative particle motion. In the following, we demonstrate that a neutrally buoyant spherical particle in a symmetry-broken Moffatt eddy flow can exhibit systematic (net) displacement across streamlines due to wall interactions, a phenomenon that has received little attention in the prior literature. These behaviors, while reminiscent of simplified steady-streaming effects (Fig. 1.3), are richer and more complex. While particles can never cross streamlines in unidirectional Stokes flow (bretherton1962motion), streamline crossing should generally be expected in the presence of wall-normal flow components. A simple example is a particle very close to a wall transported in a channel flow undergoing contraction; the particle cannot stay on its initial streamline without penetrating the wall.
@@ 3.3.1 Linear stability analysis
where the constant $c$ can vary depending on the specific flow but is ${\cal O}(1)$ (goldman1967slow2; ghalia2016sphere), cf. Sec. 2.4. Near the fixed point of the Faxen field, the $\frac{\mathcal{B}_{large}}{\mathcal{A}_{large}}$ term in $W_{y}$ dominates the others ($\frac{\mathcal{B}_{large}}{\mathcal{A}_{large}}=\mathcal{O}(a_{p}^{3}),\frac{\mathcal{C}_{large}}{\mathcal{A}_{large}}=\mathcal{O}(a_{p}^{5}),\frac{\mathcal{D}_{large}}{\mathcal{A}_{large}}=\mathcal{O}(a_{p}^{4})$) and is the only one contributing to $\mathcal{O}(a_{p}^{3})$ in (3.2). This term is still of higher $a_{p}$ order than the Faxen correction, so that, in the limits of large $\Delta$ and small $a_{p}$, the wall correction is perturbatively small.
@@ 3.5 Particle motion and manipulation in a counterclockwise vortex
Any vortex adjacent to a clockwise vortex, like the one discussed in Sec. 3.3, is counterclockwise and congruent in geometry. For example, the vortex indicated by the green frame in Fig. 2.4(a) has flow exactly reversed from the yellow-framed vortex (and a factor $\zeta$ slower). Because of the time reversibility of Stokes flow and the fact that all wall interactions result from the background Stokes flow and its derivatives, the behavior of particles on trajectories is also time-reversed. Thus, the fixed point in a counterclockwise vortex is stable, $\frac{1}{\tau}$ changes sign, and $\beta$ stays the same by definition. The limit cycle for a given $a_{p}$ is unstable but is congruent in shape with the stable cycle discussed earlier. Particles spiral inward from the unstable limit cycle towards the fixed points, but spiral outward when placed outside the limit cycle (Fig. 3.3(a)). The latter case is of prime interest because it allows particles to approach the wall more and more closely. As the gap between particle and wall diminishes, any short-ranged intermolecular attractive force (e.g. Van der Waals force) whose reach is often in the nm range (hirschfelder1954molecular; batsanov2001van) can take over and lead to attachment (sticking) of the particle to the wall. This general mechanism (hydrodynamics allows a particle to get close enough to a wall to stick by short-range attraction) has been acknowledged before (friedlander2000smoke; humphries2009filter), but the case discussed here allows for quantitative predictions.
@@
where $\kappa(x)=\frac{\partial^{2}v}{\partial y^{2}}(x,-1)$ is the background flow curvature at the wall, and the prefactor was first derived in (goren1971hydrodynamic) and confirmed in (rallabandi2017hydrodynamic). The part of the trajectory showing the exponential approach (the solid portion in Fig. 3.3(b,c)) indicates slow motion in the $x$-direction, so that $\kappa(x)$ is nearly constant. We fix the $x$-value as that of the turning point (point of maximum curvature) of the trajectory, here $x\approx 0.991$. Using this value in (3.11) gives an exponential behavior (dashed line in Fig. 3.3(c)) in good agreement with the approach of the trajectory to the wall. This behavior remains robust to changes in the exact modeling of the expansion point locations $y_{E}$ in (2.13).
@@
Our spherical-particle formalism implicitly assumes the limit of dilute particle concentration in a microfluidics application. Including the effects of particle-particle interaction is a desirable future extension of this approach, allowing for assessment of non-inertial effects in particle-laden flows, which are important for practical applications (guha2008transport).